Amount of an investment. The amount of an investment of P dollars for t years at simple interest rate r is given by A = P + Prt.
a) Rewrite this formula by factoring out the greatest common factor on the right-hand side.
b) Find A if $8300 is invested for 3 years at a simple interest rate of 15%.

Answer by checkley77(12844)

A = P + Prt.
A=P(1+rt)
A=8,300(1*3*.15)
A=8,300*.45
A=$1,710 ANS. FOR THE INTEREST.